# Hybrid quantum-classical systems: statistics, entropy, microcanonical ensemble and its connection to the canonical ensemble

### v1

@@ Abstract
We describe in detail a mathematical framework in which statistical ensembles of hybrid classical-quantum systems can be properly described. We show how a maximum entropy principle can be applied to derive the microcanonical ensemble of hybrid systems. We investigate its properties, and in particular how the microcanonical ensemble and its marginal classical and quantum ensembles can be defined for arbitrarily small range of energies for the whole system. We show how, in this situations, the ensembles are well defined for a continuum of energy values, unlike the purely quantum microcanonical ensemble, thus proving that hybrid systems translate properties of classical systems to the quantum realm. We also analyze the relation with the hybrid canonical ensemble by considering the microcanonical ensemble of a compound system composed of a hybrid subsystem weakly coupled to a reservoir and computing the marginal ensemble of the hybrid subsystem. Lastly, we apply the theory to the statistics of a toy model, which gives some insight on the different properties presented along the article.
hybrid quantum-classical systems
statistical ensembles
@@ I Introduction
The situation is however less clear in the case of hybrid quantum-classical systems. In this article, we aim to investigate statistical description and the properties of their equilibrium ensembles. Hybrid models incorporating both classical and quantum subsystems are useful in describing molecular and condensed matter systems [1, 2, 3], which present two distinctive energy scales (nuclei and electrons). In the context of field theory, hybrid quantum-classical systems have also been considered as candidates to describe quantum matter fields interacting with a (classical) gravitational field, as a semiclassical approximation or even a fundamental theory [4, 5]. Applications of hybrid models include the description of the measurement process of quantum systems, by modeling the measurement device as a classical system coupled to the quantum system to be measured [6, 7], and the study of control systems [8].
An accurate and complete description of statistical system is fundamental in many applications, such as ab initio modeling of molecules and materials and their numerical simulation methods at finite temperature [9, 10, 11, 12]. Some previous works have dealt with the description of hybrid statistical systems [13, 10, 14, 15, 16, 17], with a common approach of allowing a general and consistent description of interactions between quantum and classical ensembles. Our goal is to provide a coherent mathematical framework for the description of statistical hybrid systems, together with an appropriate definition of entropy [13] in order to present equilibrium ensembles as those maximizing the entropy of the system [18, 19, 20, 21, 22].
@@
In quantum mechanics, due to its probabilistic nature different approaches can be taken. On the one hand, statistical quantum systems can be treated in the same way as classical systems, identifying mutually exclusive events as simple events with different points of the phase space (in this case, a projective Hilbert space), with expectation values of observables as the tool to differentiate events. This approach was proposed in the mid of the 20th century [27, 28] and developed later [29, 30, 31], and it is used in fields such as quantum gravity [32, 33, 34]. However, as we discussed in a previous article [35], these models lack the extensivity properties that we expect from a thermodynamic model.
@@
The article is organized as follows. Section II summarizes the main mathematical aspects used in our description of classical, quantum and hybrid statistical systems, introducing and defining the objects encoding the state of the different types of system. Then, Section III describes the notion of hybrid entropy and how to use the maximum entropy formalism in the case of the microcanonical ensemble. In Section IV, we analyze how the microcanonical ensemble is defined in the limit of small ranges of energies of the system, proving that the marginal classical and quantum ensembles share the classical property of being defined for almost any energy, unlike the typical quantum microcanonical ensemble. In Section V we explore the relation between the microcanonical and canonical hybrid ensembles, considering the microncanonical ensemble of a compound system composed of a hybrid subsystem weakly coupled to a reservoir. Section VI discusses a simple example to obtain the microcanonical ensemble of a hybrid qubit while Section VII summarizes the main conclusions of the article and discusses potential future applications.
@@ II Mathematical description of statistical systems
In probability theory, a statistical description of a system is based on the definition of a sample space as the set of possible configurations of the system. The subsets of the sample space are called events, and a probability measure is a real function on the set of events taking values between 0 and 1 and satisfying Kolmogorov axioms [23]. In the following, we review the application of these concepts to our analysis of classical, quantum and hybrid systems.
@@ II.1 Classical states and statistics
In classical statistical systems, the sample space is identified as the phase space $M_{C}$ of the system. States $\xi\in M_{C}$ represent the values of all the variables determining the configuration of the system (e.g., positions and velocities in a mechanical system). In classical statistics, simple events are mutually exclusive: if two simple events $A_{1}=\{\xi_{1}\}$, $A_{2}=\{\xi_{2}\}$ contain different points, $\xi_{1}\neq\xi_{2}$, then conditional probabilities of $A_{1}$ with respect to $A_{2}$ and viceversa are zero.
An important element in the formalism is the description of observables. Any physical quantity that can be measured on the system (energy, temperature, position of a particle, etc.) is considered an observable, and thus is represented by a function on $M_{C}$. Mathematically, the set of classical observables $\mathcal{O}_{C}$ is a Poisson algebra of integrable smooth real valued functions on $M_{C}$, with an inner product defined by integration
@@
with $d\mu_{C}$ the volume element on $M_{C}$. Such a set can be considered (see [37]) to be the subset of real elements of the $C^{*}$–algebra of complex functions with compact domain, and the supremum norm.
@@
This definition of statistical ensembles in term of distributions allows for non-smooth functions and Dirac delta distributions to also be considered.
@@
A pure state in which the system is known to be in a certain state $\xi_{0}\in M_{C}$ is represented by the Dirac delta distribution $\delta\left(\xi-\xi_{0}\right)\in\mathcal{P}_{C}$.
@@ II.2 Quantum states and statistics
The objects presented in the description of classical statistical system are general, and an analogous approach has to be developed to describe quantum statistics. However, the first difference appears in the description of sample spaces and mutual exclusive events, which is non-trivial compared to the classical case due to the linear and probabilistic nature of quantum mechanics.
In the Schrödinger picture of quantum mechanics, the pure states of a quantum system are described by vectors in a complex Hilbert space $\mathcal{H}$. The relation between states and vectors $|\psi\rangle\in\mathcal{H}$, however, is not unique, and any two proportional non-zero vectors describe the same space. As a consequence, the space of pure states $M_{Q}$ of a quantum system can be characterized as the projective Hilbert space, which is itself diffeomorphic to the space $\mathcal{P}_{Q}^{1}$ of rank-1 projectors onto its Hilbert space $\mathcal{H}$. Using Dirac notation, the space $M_{Q}$ is described as
@@
Observables of a quantum system are represented by bounded Hermitian operators on $\mathcal{H}$, which by the Gelfand-Neimark theorem are equivalent a real unital C*-algebra $\mathcal{O}_{Q}$ [37, 38, 39]. In the case of finite-dimensional quantum systems, this is simply the algebra of Hermitian matrices of appropriate dimension. Unlike classical systems, the result of a measurement on a quantum system is not deterministic, and in general there exists uncertainty in the results of the measurement process even if the state of the system is fully determined. This is characterized by the following property: if a system is in a state $|\psi_{1}\rangle\in\mathcal{H}-\{0\}$, then the probability $p_{12}$ of finding the system in the state $|\psi_{2}\rangle\in\mathcal{H}-\{0\}$ is [40]
@@
In words, the probability of measuring a system to be in a state when we have prepared it, with absolute certainty, in a different state may be different from zero.
Due to this intrinsically probabilistic nature of quantum mechanics, a definition of a sample space as the whole set of distinct states $M_{Q}$ (as it was done in the classical case), is problematic. Given any probability measure $p$ defined on $M_{Q}$, if we consider two different simple events $\{\pi_{1}\}$ and $\{\pi_{2}\}$, they are mutually exclusive, since $\{\pi_{1}\}\cap\{\pi_{2}\}=\emptyset$ and therefore $p(\{\pi_{1}\}\cap\{\pi_{2}\})=0$. As this seems to contradict the previous statement about non-zero probabilities for different states, we must either give up the use of $M_{Q}$ as a sample space, or the interpretation of probability measures $p$ defined on it as measurement probabilities of any kind.
We will follow in this work the first approach [15, 36]. We will consider, as possible sample spaces, subsets of states in $M_{Q}$ whose associated simple events are mutually exclusive. By analogy, we will say that two states $\pi_{1}$, $\pi_{2}$ are mutually exclusive states (MESs) if ${\rm Tr}(\pi_{1}\pi_{2})=0$. It is on theses sets that we will consider statistical ensembles.
As in the case of classical systems, in a statistical description of quantum systems ensembles can be identified as linear functions on the algebra of observables. This is in fact Gleason’s theorem, which shows that any statistical ensemble is described by a density matrix, i.e., a linear function on $\mathcal{O}_{Q}$ identified as a convex combination of rank-1 projectors [39]. The corresponding set of density matrices $\mathcal{P}_{Q}$ is characterized as
$\rho^{\dagger}=\rho,\quad\rho^{2}\geq 0,\quad\mathrm{Tr}(\rho)=1,\quad\forall\rho\in\mathcal{P}_{Q}.$ (8)
@@
As indicated above, in principle it is possible to define mixed quantum states as probability distributions on the set of pure states $M_{Q}$, i.e., as a positive-definite generalized function $F_{Q}:M_{Q}\to\mathbb{R}$} normalized to one. This would imply to considering $M_{Q}$ as a sample space, analogously to the classical case. In this situation, the expectation value of an observable $\hat{A}\in\mathcal{O}_{Q}$ can be written as
@@
where the density matrix $\rho$ which corresponds to the first moment of $F_{Q}$ can be written as:
@@
Quantum dynamics is linear and therefore can be written as a master equation for the density matrix (i.e., the first moment of the distribution). We will see below that the situation is radically different for hybrid systems.
@@
Notice also that, when we choose to represent the state as a probability density over $M_{Q}$, all distributions having the same first moment, from Gleason theorem, are represented by the same density matrix. Therefore, they associate the same expectation value to any physical observable and they are hence physically indistinguishable. For example, an equiprobable distribution over the whole state space $M_{Q}$ of a 2-level quantum system is indistinguishable from a distribution of any pair of MESs with equal probability, all of which define the density matrix $\frac{1}{2}\mathbb{I}$.
@@ II.3 Hybrid classical-quantum states and statistics
Associated with that bundle, we can consider the bundle over $M_{C}$ having as fiber the $C^{*}$–algebra of quantum operators $\mathcal{O}_{Q}$. With this description, observables $\hat{\mathcal{O}}_{H}$ can be considered to correspond to the set of sections of that bundle, i.e. the set of objects of the form $\hat{A}(\xi)$, where for each $\xi\in M_{C}$, $\hat{A}(\xi)$ is supposed to define a an element of $\mathcal{O}_{Q}$, i.e., a well-defined quantum observable on $M_{Q}$. The set of hybrid operators $\mathcal{O}_{H}$ becomes isomorphic to the $C^{*}$–algebra $\mathcal{O}_{C}\otimes\mathcal{O}_{Q}$ (see [37] for the technical details concerning the definition of a suitable norm on the space of hybrid operators).
Statistical hybrid systems can be defined, by mimicking the classical and quantum cases, as probability distributions on $M_{H}$, i.e., defined by a positive-definite function $F_{QC}:M_{H}\to\mathbb{R}$, normalized to one. Expectation values are obtained, mimicking the expressions we saw above, as
$\langle\hat{A}(\xi)\rangle=\int_{M_{C}}d\mu_{C}F_{QC}(\xi,\pi)f_{A}(\xi,\pi),$ (14)
@@
$f_{A}(\xi,\pi)=\mathrm{Tr}(\hat{A}(\xi)\pi).$ (15)
As in the other cases, statistical ensembles should be described as elements in $\hat{\mathcal{O}}_{H}^{*}$. Let us provide, briefly, an appropriate context for such objects. The interested reader can find more details in [13, 37]. Description of MESs in hybrid systems can be determined following the description of MESs in purely classical and quantum systems. Two points $(\xi_{1},\pi_{1}),(\xi_{2},\pi_{2})\in M_{H}$ define MESs if and only if $\xi_{1}\neq\xi_{2}$ or $\pi_{1}\pi_{2}=\pi_{2}\pi_{1}=0$. Thus, hybrid systems share mutual-exclusivity properties with both classical and quantum systems. Let us assume that a probability distribution over $M_{H}$ is given by a probability density $F_{QC}$. If one is only interested in the classical variable and defines the corresponding marginal probability, the ensemble should behave as a classical statistical system, i.e. as a distribution in $\mathcal{P}_{C}$. On the other hand, for a fixed value of $\xi\in M_{C}$, the conditional distribution should be just quantum, thus described by a density matrix in $\mathcal{P}_{Q}$. This justifies the following description of hybrid statistical ensembles, from now called hybrid density matrices:
@@
Notice that $\rho_{Q}\neq\tilde{\rho}_{Q,\xi}$, although both are purely quantum density matrices: $\rho_{Q}$ is the quantum marginal distribution of the original hybrid distribution, while $\tilde{\rho}_{Q,\xi}$ represents the conditional distribution for a given $\xi\in M_{C}$. On the other hand, $F_{C}$ is the same marginal distribution used in (16), which imposes the following conditions on the hybrid density matrix, derived from (3):
@@
Notice that, as it was explained in [41], this hybrid density matrix represents the first quantum moment of the distribution $F_{QC}$ since

### v2

# Present address: ]Departamento de Matemáticas y Computación, Universidad de Burgos, Burgos, Spain

@@ Abstract
We describe in detail a mathematical framework in which statistical ensembles of hybrid classical-quantum systems can be properly described. We show how a maximum entropy principle can be applied to derive the microcanonical ensemble of hybrid systems. We investigate its properties, and in particular how the microcanonical ensemble and its marginal classical and quantum ensembles can be defined for arbitrarily small range of energies for the whole system. We show how, in this situation, the ensembles are well defined for a continuum of energy values, unlike the purely quantum microcanonical ensemble, thus proving that hybrid systems translate properties of classical systems to the quantum realm. We also analyze the relation with the hybrid canonical ensemble by considering the microcanonical ensemble of a compound system composed of a hybrid subsystem weakly coupled to a reservoir and computing the marginal ensemble of the hybrid subsystem. Lastly, we apply the theory to the statistics of a toy model, which gives some insight on the different properties presented along the article.
@@ I Introduction
The situation is however less clear in the case of hybrid quantum-classical systems. In this article, we aim to investigate their statistical description and the properties of their equilibrium ensembles. Hybrid models incorporating both classical and quantum subsystems are useful in describing molecular and condensed matter systems [1, 2, 3], which present two distinctive energy scales (nuclei and electrons). In the context of field theory, hybrid quantum-classical systems have also been considered as candidates to describe quantum matter fields interacting with a (classical) gravitational field, as a semiclassical approximation or even a fundamental theory [4, 5]. Applications of hybrid models include the description of the measurement process of quantum systems, by modeling the measurement device as a classical system coupled to the quantum system to be measured [6, 7, 8], and the study of control systems [9].
An accurate and complete description of statistical system is fundamental in many applications, such as ab initio modeling of molecules and materials and their numerical simulation methods at finite temperature [10, 11, 12, 13]. Some previous works have dealt with the description of hybrid statistical systems [14, 11, 15, 16, 17, 18], with a common approach of allowing a general and consistent description of interactions between quantum and classical ensembles. Choosing a consistent entropy function [14], it is possible to provide a coherent mathematical framework for the description of statistical hybrid systems which determines the hybrid canonical ensemble as the equilibrium ensemble maximizing the entropy of the system under appropriate conditions [19, 20, 21, 22, 23]. This maximization property can also be used to identify the microcanonical ensemble, as we analyze in detail in the present paper. In principle, other equilibrium ensembles, such as the grand-canonical ensemble, can also be considered within this formalism. For instance, quite recently some first-principles proposals were introduced to consider the grand-canonical ensemble for quantum systems [24, 25]. Using a similar approach, it should be possible to extend this construction to hybrid systems. That study is now in progress.
@@
In quantum mechanics, due to its probabilistic nature different approaches can be taken. On the one hand, statistical quantum systems can be treated in the same way as classical systems, identifying mutually exclusive events as simple events with different points of the phase space (in this case, a projective Hilbert space), with expectation values of observables as the tool to differentiate events. This approach was proposed in the mid of the 20th century [30, 31] and developed later [32, 33, 34], and it is used in fields such as quantum gravity [35, 36, 37]. On the other hand, as we discussed in a previous article [38], these models lack the extensivity properties that we expect from thermodynamic systems when interaction is either local or exhibits screening of large distance correlations. This limits significantly the use for systems where electromagnetic interaction plays an important role, such as molecular systems or materials.
@@
The article is organized as follows. Section II summarizes the main mathematical aspects used in our description of classical, quantum and hybrid statistical systems, introducing and defining the objects encoding the state of the different types of system. Then, Sec. III describes the notion of hybrid entropy and how to use the maximum entropy formalism in the case of the microcanonical ensemble. In Sec. IV, we analyze how the microcanonical ensemble is defined in the limit of small ranges of energies of the system, proving that the marginal classical and quantum ensembles share the classical property of being defined for almost any energy, unlike the typical quantum microcanonical ensemble. In Sec. V we explore the relation between the microcanonical and canonical hybrid ensembles, considering the microcanonical ensemble of a compound system composed of a hybrid subsystem weakly coupled to a reservoir. Section VI discusses a simple example to obtain the microcanonical ensemble of a hybrid qubit while Sec. VII summarizes the main conclusions of the article and discusses potential future applications.
@@ II Mathematical description of statistical systems
In probability theory, a statistical description of a system is based on the definition of a sample space as the set of possible configurations of the system. The subsets of the sample space are called events, and a probability measure is a real function on the set of events taking values between 0 and 1 and satisfying Kolmogorov axioms [26]. This is achieved by considering a probability distribution, which involves a definition of a sample space and a probability measure defined on it. This is trivial to do in the classical world, but the quantum case is more subtle. In the following, we review the application of these concepts to our analysis of classical, quantum and hybrid systems.
@@ II.1 Classical states and statistics
In classical statistical systems, the sample space is identified with the phase space $M_{C}$ of the system. States $\xi\in M_{C}$ represent the values of all the variables determining the configuration of the system (e.g., positions and velocities in a mechanical system). In classical statistics, simple events are mutually exclusive: if two simple events $A_{1}=\{\xi_{1}\}$, $A_{2}=\{\xi_{2}\}$ contain different points, $\xi_{1}\neq\xi_{2}$, then conditional probabilities of $A_{1}$ with respect to $A_{2}$ and viceversa are zero, i.e., the probability of detecting the system in state $\xi_{1}$ after it has been prepared in $\xi_{2}$ is zero, and viceversa.
An important element in the formalism is the description of observables. Any physical quantity that can be measured on the system (energy, temperature, position of a particle, etc.) is considered an observable, and thus is represented by a function on $M_{C}$. Mathematically, the set of classical observables $\mathcal{O}_{C}$ is a Poisson algebra [^1] This requires the existence of a Poisson bracket on the algebra of functions, which allows us to define Hamiltonian dynamics. The Poisson bracket is just a skew-symmetric bidifferential operator that satisfies the Jacobi identity when acting on three arbitrary functions. See Ref. [40] for details. of integrable smooth real valued functions on $M_{C}$, with an inner product defined by integration
@@
with $d\mu_{C}$ the volume element on $M_{C}$. Such a set can be considered (see [41]) to be the subset of real elements of the $C^{*}$ algebra of complex functions with compact domain, and the supremum norm:
@@
This definition of statistical ensembles in term of distributions allows for nonsmooth functions and Dirac $\delta$ distributions to also be considered.
@@
A pure state in which the system is known to be in a certain state $\xi_{0}\in M_{C}$ is represented by the Dirac $\delta$ distribution $\delta\left(\xi-\xi_{0}\right)\in\mathcal{P}_{C}$.
@@ II.2 Quantum states and statistics
The objects presented in the description of classical statistical system are general, and an analogous approach has to be developed to describe quantum statistics. However, the first difference appears in the description of sample spaces and mutual exclusive events, which is nontrivial compared to the classical case due to the linear and probabilistic nature of quantum mechanics.
In the Schrödinger picture of quantum mechanics, the pure states of a quantum system are described by vectors in a complex Hilbert space $\mathcal{H}$. The relation between states and vectors $|\psi\rangle\in\mathcal{H}$, however, is not unique, and any two proportional nonzero vectors describe the same state. As a consequence, the space of pure states $M_{Q}$ of a quantum system can be characterized as the projective Hilbert space, which is itself diffeomorphic to the space $\mathcal{P}_{Q}^{1}$ of rank-1 projectors onto its Hilbert space $\mathcal{H}$. Using Dirac notation, the space $M_{Q}$ is described as
@@
Observables of a quantum system are represented by self-adjoint operators on $\mathcal{H}$, which constitute the real (with respect to the canonical involution) part of a unital $C^{*}$ algebra $\mathcal{O}_{Q}$ [41, 42, 43]. In the case of finite-dimensional quantum systems, this set corresponds simply to the algebra of Hermitian matrices of appropriate dimension. Unlike classical systems, the result of a measurement on a quantum system is not deterministic, and in general there exists uncertainty in the results of the measurement process even if the state of the system is fully determined. This is characterized by the following property: if a system is in a state $|\psi_{1}\rangle\in\mathcal{H}-\{0\}$, then the probability $p_{12}$ of finding the system in the state $|\psi_{2}\rangle\in\mathcal{H}-\{0\}$ is [44]
@@
In words, the probability of measuring a system to be detected in a state when we have prepared it, with absolute certainty, in a different state may be different from zero.
Due to this intrinsically probabilistic nature of quantum mechanics, a definition of a sample space as the whole set of distinct states $M_{Q}$ (as it was done in the classical case), is problematic. Given any probability measure $p$ defined on $M_{Q}$, if we consider two different simple events $\{\pi_{1}\}$ and $\{\pi_{2}\}$, they are mutually exclusive, since $\{\pi_{1}\}\cap\{\pi_{2}\}=\emptyset$ and therefore $p(\{\pi_{1}\}\cap\{\pi_{2}\})=0$. This seems to contradict the previous statement about nonzero probabilities for different states, but it does not: a probability measure defined on $M_{Q}$ must be interpreted as a preparation measure. Hence, $p(\pi_{1})$ is the probability that the system has been prepared in state $\pi_{1}$. If $p(\pi_{1})=1$, then there is obviously zero probability that it has been prepared in a different state $\pi_{2}$. But then, as shown in Eq. (7), there is a nonzero probability of being detected or measured in that different state $\pi_{2}$, if it is not orthogonal to $\pi_{1}$.
In this work, our approach will be the following [16, 39]: we will say that two states $\pi_{1}$, $\pi_{2}$ are mutually exclusive states (MESs) if ${\rm Tr}(\pi_{1}\pi_{2})=0$. Thus, we are considering a measuring probability. To properly define it, one must consider, as possible sample spaces, subsets of states in $M_{Q}$ whose associated simple events are mutually exclusive in that definition.
How to represent a statistical state of the quantum system, then? The state of this statistical system can be always represented by a density matrix of the given system. This is in fact the consequence of Gleason’s theorem [45]: any measure of a quantum system (with Hilbert space’s dimension larger than 2) can be written as expectation values with respect to a unique self-adjoint, positive and trace-class operator, with trace equal to one. This defines a density matrix of said system, which in this formalism can be identified with a linear functional on $\mathcal{O}_{Q}$ and written as a convex combination of rank-1 projectors [43]. The corresponding set of density matrices $\mathcal{P}_{Q}$ contains those linear operators $\rho$ on $\mathcal{H}$ such that
$\rho^{\dagger}=\rho,\quad\rho\geq 0,\quad\mathrm{Tr}(\rho)=1.$ (8)
@@
Thus we see that the state of the system determines a linear functional on the $C^{*}$ algebra of observables, i.e. an element of the dual space $\mathcal{O}_{Q}^{*}$.
In principle it is also possible to define mixed quantum states as probability distributions on the set of pure states $M_{Q}$, i.e., as a positive-definite generalized function $F_{Q}:M_{Q}\to\mathbb{R}$} normalized to one. This implies considering $M_{Q}$ as a sample space, analogously to the classical case. In this situation, the expectation value of an observable $\hat{A}\in\mathcal{O}_{Q}$ can be written as
@@
where the density matrix $\rho$ which corresponds to the first moment of $F_{Q}$ can be written as
@@
In this way, independence of the different events is captured in the eigenspaces of the density matrix. Nonetheless, notice that two different probability distributions $F_{Q}$ are physically equivalent as long as they have the same first moment. No measurement can distinguish between the two distributions and hence they must correspond to the same physical state.
Quantum dynamics is linear on the space of states and therefore can be written as a master equation for the density matrix (i.e., the first moment of the distribution). We will see below that the situation is radically different for hybrid systems.
@@
Notice also that, when we choose to represent the state as a probability density over $M_{Q}$, all distributions having the same first moment, from Gleason theorem, are represented by the same density matrix. Therefore, they associate the same expectation value to any physical observable and they are hence physically indistinguishable. For example, an equiprobable distribution over the whole state space $M_{Q}$ of a two-level quantum system is indistinguishable from a distribution of any pair of MESs with equal probability, all of which define the density matrix $\frac{1}{2}\mathbb{I}$.
@@ II.3 Hybrid classical-quantum states and statistics
Associated with that bundle, we can consider the bundle over $M_{C}$ having as fiber the $C^{*}$ algebra of quantum operators $\mathcal{O}_{Q}$. With this description, observables $\hat{\mathcal{O}}_{H}$ can be considered to correspond to the set of sections of that bundle, i.e. the set of objects of the form $\hat{A}(\xi)$, where for each $\xi\in M_{C}$, $\hat{A}(\xi)$ defines an element of $\mathcal{O}_{Q}$, i.e., a well-defined quantum observable on $M_{Q}$. The set of hybrid operators $\mathcal{O}_{H}$ becomes isomorphic to the $C^{*}$ algebra $\mathcal{O}_{C}\otimes\mathcal{O}_{Q}$ (see Ref. [41] for the technical details concerning the definition of a suitable norm on the space of hybrid operators).
Statistical hybrid systems can be defined, by mimicking the classical and quantum cases, as probability distributions on $M_{H}$, i.e., defined by a positive-definite function $F_{QC}:M_{H}\to\mathbb{R}$, normalized to one. Expectation values are obtained as
$\langle\hat{A}(\xi)\rangle=\int_{M_{H}}d\mu_{H}F_{QC}(\xi,\pi)f_{A}(\xi,\pi),$ (14)
@@
$f_{A}(\xi,\pi)=\mathrm{Tr}(\hat{A}(\xi)\pi),\quad d\mu_{H}=d\mu_{C}d\mu_{Q}.$ (15)
As in the other cases, statistical ensembles should be described as elements in $\hat{\mathcal{O}}_{H}^{*}$. Let us provide, briefly, an appropriate context for such objects. The interested reader can find more details in Refs. [14, 41]. Description of MESs in hybrid systems can be determined following the description of MESs in purely classical and quantum systems. Two points $(\xi_{1},\pi_{1}),(\xi_{2},\pi_{2})\in M_{H}$ define MESs if and only if $\xi_{1}\neq\xi_{2}$ or $\pi_{1}\pi_{2}=\pi_{2}\pi_{1}=0$. Thus, hybrid systems share mutual-exclusivity properties with both classical and quantum systems. Let us assume that a probability distribution over $M_{H}$ is given by a probability density $F_{QC}$. If one is only interested in the classical variable and defines the corresponding marginal probability, then the ensemble should behave as a classical statistical system, i.e. as a distribution in $\mathcal{P}_{C}$. However, for a fixed value of $\xi\in M_{C}$, the conditional distribution should be just quantum, and hence from Gleason theorem described by a density matrix in ${P}_{Q}$. This justifies the following description of hybrid statistical ensembles, from now called hybrid density matrices:
@@
Notice that $\rho_{Q}\neq\tilde{\rho}_{Q,\xi}$, although both are purely quantum density matrices: $\rho_{Q}$ is the quantum marginal distribution of the original hybrid distribution, while $\tilde{\rho}_{Q,\xi}$ represents the conditional distribution for a given $\xi\in M_{C}$. However, $F_{C}$ is the same marginal distribution used in Eq. (16), which imposes the following conditions on the hybrid density matrix, derived from Eq. (3):
@@
Notice that, as it was explained in Ref. [48], this hybrid density matrix represents the first quantum moment of the distribution $F_{QC}$ since